Calculate the mass that would yeild a body mass index of 20 or less for the following heights
a)150 cm
b) 185 cm
An altheletes BMI is 20. For each mass, how tall is the athele?
a)65 kg
b)55 kg

Answer by nerdybill(5412)

Calculate the mass that would yeild a body mass index of 20 or less for the following heights
a)150 cm
b) 185 cm
An altheletes BMI is 20. For each mass, how tall is the athele?
a)65 kg
b)55 kg
.
First, you need to know the formula to calculate the "body mass index":
BMI = weight(kg)/height^2(m^2)
where
BMI is the body mass index
.
a)150 cm
100 centimeter = 1 meter
converting to meters:
150 cm = 1.5m
Pluggin in our values into:
BMI = weight(kg)/height^2(m^2)
20 = w/(1.5)^2
w = 20(1.5)^2
w = 45 kg
.
b) 185 cm
100 centimeter = 1 meter
converting to meters:
185 cm = 1.85m
Pluggin in our values into:
BMI = weight(kg)/height^2(m^2)
20 = w/(1.85)^2
w = 20(1.85)^2
w = 68.45 kg
.
An altheletes BMI is 20. For each mass, how tall is the athele?
a)65 kg
Pluggin in our values into:
BMI = weight(kg)/height^2(m^2)
20 = 65/h^2
h^2 = 65/20
h^2 = 3.25
h = 1.8 m or 180 cm
.
b)55 kg
Pluggin in our values into:
BMI = weight(kg)/height^2(m^2)
20 = 55/h^2
h^2 = 55/20
h^2 = 2.75
h = 1.66 m or 166 cm